# On the Hidden Mystery of OCR in Large Multimodal Models

### v5

@@ Abstract
Large models have recently played a dominant role in natural language processing and multimodal vision-language learning. However, their effectiveness in text-related visual tasks remains relatively unexplored. In this paper, we conducted a comprehensive evaluation of Large Multimodal Models, such as GPT4V and Gemini, in various text-related visual tasks including Text Recognition, Scene Text-Centric Visual Question Answering (VQA), Document-Oriented VQA, Key Information Extraction (KIE), and Handwritten Mathematical Expression Recognition (HMER). To facilitate the assessment of Optical Character Recognition (OCR) capabilities in Large Multimodal Models, we propose OCRBench, a comprehensive evaluation benchmark. Our study encompasses 29 datasets, making it the most comprehensive OCR evaluation benchmark available. Furthermore, our study reveals both the strengths and weaknesses of these models, particularly in handling multilingual text, handwritten text, non-semantic text, and mathematical expression recognition. Most importantly, the baseline results showcased in this study could provide a foundational framework for the conception and assessment of innovative strategies targeted at enhancing zero-shot multimodal techniques. The evaluation pipeline and benchmark are available at https://github.com/Yuliang-Liu/MultimodalOCR.
@@ 1 Introduction
To this end, we undertake a comprehensive study on 14 LMMs by evaluating their OCR ability on five representative tasks: text recognition, Scene Text-Centric VQA, Document-Oriented VQA, key information extraction and handwritten mathematical expression recognition. Our findings indicate that even state-of-the-art large multimodal models, such as  [16] and GPT4V [17], encounter difficulties in recognizing handwritten mathematical expressions. LMMs continue to struggle with blurry text images, handwritten text, multilingual text. Moreover, they heavily rely on semantic understanding to recognize words, often favoring common words over random letter sequences. Consequently, these findings highlight even the most powerful LMM leaves a gap to match the domain-specific methods in various traditional text tasks, and even worse for more complicated tasks. However, it highlights the vast opportunity to improve the OCR ability of LMM by adapting them to the OCR domains.
@@ 2.1 Large Multimodal Models
The remarkable success of Large Language Models (LLMs) has paved the way for the development of large multimodal models, which combine pretrained visual models with LLMs to enable their visual capabilities. BLIP2 [18] introduces the Querying Transformer (Q-Former) as a means to bridge the gap between vision and language models. Flamingo [13] and OpenFlamingo [19] enhance a frozen pretrained LLM by incorporating novel gated cross-attention-dense layers, enabling conditioning on visual inputs. LLaVA [15] pioneers the use of GPT-4 to generate multimodal instruction-following data. Other works, such as [20, 21], also focus on aligning the vision module and LLM for improved multimodal understanding. Additionally, [22, 23] emphasize modality collaboration for image and text. LLaVAR [24] collects training data with rich text and uses a higher-resolution CLIP as the visual encoder to enhance LLaVA’s OCR ability. BLIVA [25] combines instruction-aware and global visual features to capture richer image information. MiniGPT4V2 [26] uses unique identifiers for different tasks when training the model to better distinguish each task instruction effortlessly. UniDoc [27] performs unified multimodal instruction tuning on large-scale instruction following datasets and leverages the beneficial interactions among tasks to enhance the performance of each individual task. Docpedia [28] directly processes visual input in the frequency domain rather than the pixel space. Monkey [29] enhanced the LMM’s ability to perceive details at a low cost by the generated detailed caption and a high-resolution model architecture.
@@ 2.2 Benchmarks for Large Multimodal Models
With the ongoing advancements in Large Multimodal Models, the question of how to effectively evaluate their performance has emerged. [15, 30, 22] introduces GPT-4 or human evaluation to assess the output of LMMs. MMBench [31] evalute LMM using multiple-choice questions across various dimensions of ability. MME [32] measures both perception and cognition abilities on True or False question. However, their testing on OCR data is limited. Additionally, the True/False question or multiple-choice question cannot accurately assess the OCR ability to recognize words in an image. In this paper, we conducted an extensive study on various LMMs for five prominent text-related tasks. To facilitate the evaluation of LMMs’ OCR ability, we also present OCRBench, a collection of 1000 manually filtered and corrected question-answer pairs on five representative text-related tasks.
@@ 3.1 Evaluation Metric and Evaluation Dataset
Text Recognition: We evaluate LMM using widely-adopted OCR text recognition datasets, including (1) Regular Text Recognition: IIIT5K [34], SVT [35], IC13 [36]; (2)Irregular Text Recognition: IC15 [37], SVTP [38], CT80 [39], COCOText(COCO) [40], SCUT-CTW1500 (CTW) [41], Total-Text (TT) [42]; (3) Occlusion Scene Text [43], encompassing weakly occluded scene text (WOST) and heavily occluded scene text (HOST); (4) Artistic Text Recognition: WordArt [44]; (5) Handwirtten Text Recognition: IAM [45]; (6) Chinese Text Recognition: ReCTS [46]; (7) Handwritten Digit String Recognition: ORAND-CAR-2014(CAR-A) [47]; (8) Non-Semantic Text(NST) and Semantic Text(ST): LMMs primarily rely on semantic understanding to recognize words. In the experiments, we find that the LMMs have poor recognition performance on character combinations that lacked semantics. To confirm this, we create two datasets: Semantic Text (ST) and Non-Semantic Text (NST) using the IIIT5k dictionary. The ST dataset consists of 3000 images with words from the IIIT5K dictionary, while the NST dataset contains the same words but with shuffled characters without semantics. For english words, we employ the consistent prompt "what is written in the image?". For Chinese words in ReCTS dataset, we adapt the prompt to "What are the Chinese characters in the image?". For handwritten digit string, we utilize the prompt "what is the number in the image?".
@@
Key Information Extraction: We conduct experiments on SROIE [55], FUNSD [56] and POIE [57]. SROIE contains 1000 complete scanned receipt images for OCR and key information extraction competitions. In this competition, it is required to extract the company, date, address, and total expenditure information based on the receipts. FUNSD dataset consists of 199 real, fully annotated, scanned forms that may contain noise. POIE consists of camera images from Nutrition Facts label of products in English and 3,000 images with 111,155 text instances are collected. KIE dataset requires to extract key-value pairs in the image. To enable LMMs to accurately extract the correct value for a given key in the KIE dataset, we employ manual prompt design. For SROIE dataset, we utilize the following prompts to assist LMMs in generating the respective values for "company", "date", "address", and "total": "what is the name of the company that issued this receipt?", "when was this receipt issued?", "where was this receipt issued?", and "what is the total amount of this receipt?". Additionally, to retrieve the corresponding value for a given key in FUNSD and POIE, we utilize the prompt "What is the value for ‘{key}’?"
Handwritten Mathematical Expression Recognition(HMER): We evaluate on HME100K [58], which consists of 74,502 images for training and 24,607 images for testing, with 245 symbol classes. During evaluation, we use the prompt "Please write out the expression of the formula in the image using LaTeX format.".
@@ 3.2 Results
The results of text recognition are shown in Tab. 1. LMMs achieve comparable performance to state-of-the-art supervised models in recognizing regular text, irregular text, occluded text, and artistic text. Even in the WordArt dataset, which primarily features challenging artistic text, InstructBLIP2 and BLIVA surpass the performance of supervised state-of-the-art model. However, LMMs perform poorly in recognizing handwritten text, Chinese text, handwritten strings, and non-semantic text. In tasks such as Scene Text-Centric VQA, Document-Oriented VQA, and KIE, LMMs with smaller input resolutions consistently yield poorer results compared to those with larger input sizes. This is because texts often contain intricate structures and can vary in size, requiring the LMMs to capture fine details. The results are shown in Tab. 2 LMMs struggle to complete the complex task of recognizing handwritten mathematical expressions. Through extensively analyzing the results, we qualitatively summarize the limitations of LMMs in text tasks:
  • Semantic-Reliance. LMMs primarily rely on semantic understanding to recognize words. In our experiments, we observed that LMMs exhibit poor recognition performance when it comes to character combinations that lack semantic meaning. Specifically, when we altered the order of letters in each word, the accuracy of LMMs on the NST dataset decreased by an average of 57.0% compared to the ST dataset, while the SOTA method for scene text recognition only drops by around 4.6%. We believe this is because the SOTA method for scene text recognition directly recognizes each character, and semantic information is just used to assist the recognition process, while LMMs primarily rely on semantic understanding to recognize words. This finding is further supported by the low accuracy observed on the ORAND dataset. As shown in Fig. 1, LMM successfully identified the word "Message," but incorrectly recognized "egaesMs," which is a rearranged version of the word "Message."
  • Handwritten Text. LMMs may encounter difficulties in accurately recognizing handwritten text due to various reasons, including the resemblances in shapes between handwritten letters and numbers. Handwritten text often appears incomplete or blurry due to factors like fast writing speed, irregular handwriting, or low-quality paper. On average, LMMs perform 51.9% worse than supervised state-of-the-art model in this task.
  • Multilingual Text. The notable performance gap observed in LMMs on ReCTS, ESTVQA(En), and ESTVQA(Ch) highlights their limited proficiency in the Chinese language, as indicated in Tab. 1 and Tab. 2. Among the LMMs, accurately recognizing Chinese words or responding to Chinese questions proves to be a challenge. Monkey outperforms other LMMs in Chinese scenarios due to its LLM and visual encoder being trained on a substantial amount of Chinese data.
  • Fine-grain Perception. Currently, most LMMs have their input images limited to a resolution of 224 x 224, consistent with the input size of the visual encoder used in their architecture. However, high-resolution input images can capture more details from the image. Due to the limited input resolution of LMMs such as BLIP2, their capacity to extract fine-grained information in tasks like Scene Text-Centric VQA, Document-Oriented VQA, and KIE is constrained. However, LMMs such as Monkey, which can support a resolution of 1344×896, exhibit improved performance in these specific tasks.
  • HMER. LMMs struggle to recognize handwritten mathematical expressions due to the presence of messy handwritten characters, complex spatial structures, and the indirect LaTeX representation. The lack of training data for the handwritten mathematical expression recognition task is also a crucial factor.
@@ 3.3 OCRBench
Evaluating multiple datasets can be time-consuming, and the inaccurate annotations in some datasets make accuracy-based evaluation less precise. Realizing these limitations, we have developed OCRBench to facilitate the accurate and convenient evaluation of LMMs’ OCR capabilities. OCRBench consists of five components: text recognition, Scene Text-Centric VQA, Document-Oriented VQA, KIE, and HMER. It includes 1000 question-answer pairs, and for the KIE task, we added the prompt "Answer this question using the text in the image directly." to restrict the model’s response format. The specific composition of OCRBench is shown in appendix. To ensure a more accurate evaluation, we manually verified and corrected incorrect answers for the 1000 question-answer pairs, providing alternative correct answer candidates. The final results are presented in the Tab. 3, with Gemini achieving the highest score and GPT4V securing the second position. It is important to note that due to rigorous safety reviews by OPEN AI, GPT4V refused to provide results for 84 images in OCRBench. Monkey demonstrated OCR capabilities that were only inferior to GPT4V and Gemini. From the test results, we can observe that even state-of-the-art models like GPT4V and Gemini struggle with the HMER task. Furthermore, they also face challenges in handling unclear images, handwritten text, Non-Semantic text and adhering to task instructions. As shown in Fig. 2 (g), even when explicitly requested to answer using the text found in the image, Gemini consistently interprets "02/02/2018" as "2 February 2018".
Figure 2: Some erroneous results from Gemini and GPT-4V. ‘[Q]’ represents question, ‘[G]’ represents ground truth. (a) showcases results on blurry image, (b) presents results on handwritten text, (c) and (d) showcase results on Non-Semantic text, (e) and (f) showcase results on HMER, (g) displays instances where the task instructions are not adhered to.
4 Discussions
@@
Why LMMs work for OCR? While we offer some analysis based on the results, the question as to why these multimodal models can deliver acceptable performance on OCR tasks remains challenging to conclusively explain. One plausible explanation for the success of multimodal models in Optical Character Recognition (OCR) tasks lies in the training of multimodal models (similar to CLIP), which we believe includes some OCR data. However, unsupervised text-image pairs in the training data cannot compete with fully supervised data.
@@
From a network perspective, the pre-trained models of the visual encoder and Large Language Model (LLM) already demonstrate a strong grasp of their respective domain data, each operating within their structured feature space. These LMMs bridge visual and language information using components such as a linear projection layer, that functions as a step of visual tokenization. This process aligns visual tokens within the pre-trained language model’s word embedding space, resulting in visual embeddings that closely mirror their corresponding word embeddings. This alignment facilitates text recognition, and the LLM subsequently presents this OCR data to users in a generative manner.
Future research could benefit from an ablation analysis to understand the impact of the volume of multimodal training data on OCR performance.
Future Directions: Despite demonstrating promise, current LMMs struggle with complex tasks and cannot rival domain-specific methods in traditional text tasks. To improve, these models should focus on refining their capabilities in detecting fine-grained features in images, enhancing their understanding of individual character shapes, and building multilingual data for training.

### v6

@@ Abstract
Large models have recently played a dominant role in natural language processing and multimodal vision-language learning. However, their effectiveness in text-related visual tasks remains relatively unexplored. In this paper, we conducted a comprehensive evaluation of Large Multimodal Models, such as GPT4V and Gemini, in various text-related visual tasks including Text Recognition, Scene Text-Centric Visual Question Answering (VQA), Document-Oriented VQA, Key Information Extraction (KIE), and Handwritten Mathematical Expression Recognition (HMER). To facilitate the assessment of Optical Character Recognition (OCR) capabilities in Large Multimodal Models, we propose OCRBench, a comprehensive evaluation benchmark. OCRBench contains 29 datasets, making it the most comprehensive OCR evaluation benchmark available. Furthermore, our study reveals both the strengths and weaknesses of these models, particularly in handling multilingual text, handwritten text, non-semantic text, and mathematical expression recognition. Most importantly, the baseline results presented in this study could provide a foundational framework for the conception and assessment of innovative strategies targeted at enhancing zero-shot multimodal techniques. The evaluation pipeline and benchmark are available at https://github.com/Yuliang-Liu/MultimodalOCR.
@@ 1 Introduction
To this end, we conduct a comprehensive study on 14 LMMs by evaluating their OCR ability on five representative tasks: Text Recognition, Scene Text-Centric VQA, Document-Oriented VQA, Key Information Extraction, and Handwritten Mathematical Expression Recognition. The results indicate that even state-of-the-art large multimodal models, such as Gemini [16] and GPT4V [17], still encounter challenges in recognizing blurry text images, handwritten text, multilingual text, and handwritten mathematical expressions. Moreover, we observe that they heavily rely on semantic understanding to recognize words, often favoring common words over random letter sequences. The above findings demonstrate that even the most powerful LMM still exhibits significant gaps compared to the domain-specific methods in various text-related tasks. Consequently, there exists a promising opportunity to enhance the OCR capabilities of LMMs through domain-specific adaptations and optimizations.
@@ 2.1 Large Multimodal Models
The remarkable success of Large Language Models (LLMs) has paved the way for the development of large multimodal models, which combine pretrained visual models with LLMs to enable their visual capabilities. BLIP2 [18] introduces the Querying Transformer (Q-Former) as a means to bridge the gap between vision and language models. Flamingo [13] and OpenFlamingo [19] enhance a frozen pretrained LLM by incorporating novel gated cross-attention-dense layers, enabling conditioning on visual inputs. LLaVA [15] pioneers the use of GPT-4 to generate multimodal instruction-following data. Other works, such as [20, 21], also focus on aligning the vision module and LLM for improved multimodal understanding. Additionally, [22, 23] emphasize modality collaboration for image and text. LLaVAR [24] collects training data with rich text and uses a higher-resolution CLIP as the visual encoder to enhance LLaVA’s OCR ability. BLIVA [25] combines instruction-aware and global visual features to capture richer image information. MiniGPT4V2 [26] uses unique identifiers for different tasks when training the model to better distinguish each task instruction effortlessly. UniDoc [27] performs unified multimodal instruction tuning on large-scale instruction following datasets and leverages the beneficial interactions among tasks to enhance the performance of each individual task. Docpedia [28] directly processes visual input in the frequency domain rather than the pixel space. Monkey [29] enhanced the LMM’s ability to perceive details at a low cost by the generated detailed caption and a high-resolution model architecture. TextMonkey[30] introduced window attention to strengthen the correlation between different patches and introduced a token resampler to reduce the length of image tokens.
@@ 2.2 Benchmarks for Large Multimodal Models
With the ongoing advancements in Large Multimodal Models, the question of how to effectively evaluate their performance has emerged. [31, 32] have developed effective systems to assess LMMs. [15, 33, 22] introduces GPT-4 or human evaluation to assess the output of LMMs. MMBench [34] evaluates LMM using multiple-choice questions across various dimensions of ability. MME [35] measures both perception and cognition abilities on True or False questions. However, their testing on OCR data is limited. Additionally, the True/False question or multiple-choice question cannot accurately assess the OCR’s ability to recognize words in an image. In this paper, we conducted an extensive study on various LMMs for five prominent text-related tasks. To facilitate the evaluation of LMMs’ OCR ability, we also present OCRBench, a collection of 1000 manually filtered and corrected question-answer pairs on five representative text-related tasks.
@@ 3.1 Evaluation Metric and Evaluation Dataset
Text Recognition: We evaluate LMM using widely-adopted OCR text recognition datasets, including (1) Regular Text Recognition: IIIT5K [37], SVT [38], IC13 [39]; (2)Irregular Text Recognition: IC15 [40], SVTP [41], CT80 [42], COCOText(COCO) [43], SCUT-CTW1500 (CTW) [44], Total-Text (TT) [45]; (3) Occlusion Scene Text [46], encompassing weakly occluded scene text (WOST) and heavily occluded scene text (HOST); (4) Artistic Text Recognition: WordArt [47]; (5) Handwirtten Text Recognition: IAM [48]; (6) Chinese Text Recognition: ReCTS [49]; (7) Handwritten Digit String Recognition: ORAND-CAR-2014(CAR-A) [50]; (8) Non-Semantic Text(NST) and Semantic Text(ST): LMMs primarily rely on semantic understanding to recognize words. In the experiments, we found that the LMMs have poor recognition performance on character combinations that lacked semantics. To confirm this, we create two datasets: Semantic Text (ST) and Non-Semantic Text (NST) using the IIIT5k dictionary. The ST dataset consists of 3000 images with words from the IIIT5K dictionary, while the NST dataset contains the same words but with shuffled characters without semantics. For English words, we employ the consistent prompt “What is written in the image”. For Chinese words in ReCTS dataset, we adapt the prompt to “What are the Chinese characters in the image”. For handwritten digit strings, we utilize the prompt “What is the number in the image?”.
@@
Key Information Extraction: We conduct experiments on SROIE [58], FUNSD [59] and POIE [60]. SROIE contains 1000 complete scanned receipt images for OCR and key information extraction competitions. In this competition, the company, date, address, and total expenditure information must be extracted based on the receipts. FUNSD dataset consists of 199 real, fully annotated, scanned forms that may contain noise. POIE consists of camera images from the Nutrition Facts label of products in English and 3,000 images with 111,155 text instances are collected. The KIE dataset requires the extraction of key-value pairs in the image. To enable LMMs to accurately extract the correct value for a given key in the KIE dataset, we employ manual prompt design. For the SROIE dataset, we utilize the following prompts to assist LMMs in generating the respective values for “company”, “date”, “address”, and “total”: “What is the name of the company that issued this receipt?”, “When was this receipt issued?”, “Where was this receipt issued?”, and “What is the total amount of this receipt?”. Additionally, to retrieve the corresponding value for a given key in FUNSD and POIE, we utilize the prompt “What is the value for ‘{key}’?”
@@
Handwritten Mathematical Expression Recognition(HMER): We evaluate on HME100K [61], which consists of 74,502 images for training and 24,607 images for testing, with 245 symbol classes. During evaluation, we use the prompt “Please write out the expression of the formula in the image using LaTeX format.”.
@@ 3.2 Results
The results of text recognition are shown in Tab. 1. LMMs achieve comparable performance to state-of-the-art supervised models in recognizing regular text, irregular text, occluded text, and artistic text. Particularly in the WordArt dataset, which predominantly comprises challenging artistic text, InstructBLIP2, and BLIVA even outperform the supervised state-of-the-art model. However, LMMs exhibit poor performance in recognizing handwritten text, Chinese text, handwritten strings, and non-semantic text. In tasks such as Scene Text-Centric VQA, Document-Oriented VQA, and KIE, LMMs with smaller input resolutions consistently produce poorer results compared to those with larger input sizes. This is due to the intricate structures and varying sizes of texts, requiring LMMs to capture fine details. The results are shown in Tab. 2. LMMs face challenges in accurately recognizing handwritten mathematical expressions. By extensively analyzing the results, we provide a qualitative overview of the limitations of LMMs in text-related tasks:
  • Semantic-Reliance. LMMs primarily rely on semantic understanding to recognize words. In our experiments, we observed that LMMs exhibit poor recognition performance when it comes to character combinations that lack semantic meaning. Specifically, when we altered the order of letters in each word, the accuracy of LMMs on the NST dataset decreased by an average of 57.0% compared to the ST dataset, while the SOTA method for scene text recognition only drops by around 4.6%. We believe this is because the SOTA method for scene text recognition directly recognizes each character, and semantic information is just used to assist the recognition process, while LMMs primarily rely on semantic understanding to recognize words. This finding is further supported by the low accuracy observed on the ORAND dataset. As shown in Fig. 1, LMM successfully identified the word “Message”, but incorrectly recognized “egaesMs”, which is a rearranged version of the word “Message”.
@@
  • Handwritten Text. LMMs may encounter difficulties in accurately recognizing handwritten text due to various reasons, including the resemblances in shapes between handwritten letters and numbers. Handwritten text often appears incomplete or blurry due to factors like fast writing speed, irregular handwriting, or low-quality paper. On average, LMMs exhibit a 51.9% lower performance compared to the supervised state-of-the-art model in this particular task.
  • Multilingual Text. As indicated in Tab. 1 and Tab. 2, there exists a notable performance gap between ESTVQA(CN) and ESTVQA(EN). The LMMs achieve limited proficiency in the Chinese language. Accurately recognizing Chinese words or responding to Chinese questions poses a considerable challenge for LMMs. Training LMMs with Chinese data emerges as a viable solution. For example, Monkey surpasses other LMMs in Chinese scenarios due to the substantial training of its LLM and visual encoder on Chinese data.
  • Fine-grain Perception. Currently, the resolution of most LMMs is limited to 224 x 224 due to the visual encoder used in their architecture. However, supporting high-resolution input is essential for LMMs to capture finer details within images. The restricted input resolution of LMMs like BLIP2 hinders their ability to extract detailed information in tasks such as Scene Text-Centric VQA, Document-Oriented VQA, and KIE. Conversely, LMMs like Monkey, which can handle a resolution of 1344×896, demonstrate enhanced performance in these specific tasks.
  • HMER. Recognizing handwritten mathematical expressions poses a challenge for LMMs due to the presence of messy handwritten characters, complex spatial structures, and the indirect LaTeX representation. Additionally, the scarcity of training data for this specific task further complicates the recognition process for LMMs.
Figure 2: Some erroneous results from Gemini and GPT-4V. ‘[Q]’ represents question, ‘[G]’ represents ground truth. (a) presents the results on blurry image, (b) presents the results on handwritten text, (c) and (d) present the results on Non-Semantic text, (e) and (f) present the results on HMER, (g) display instances where the task instructions are not adhered to.
@@ 3.3 OCRBench
Evaluating the model on multiple datasets is time-consuming, and the presence of inaccurate annotations in some datasets diminishes the precision of accuracy-based evaluations. To solve these issues, we develop OCRBench to facilitate the accurate and convenient evaluation of LMMs’ OCR capabilities. OCRBench consists of five components: text recognition, Scene Text-Centric VQA, Document-Oriented VQA, KIE, and HMER. It includes 1000 question-answer pairs, and for the KIE task, we added the prompt “Answer this question using the text in the image directly.” to restrict the model’s response format. The specific composition of OCRBench is shown in the appendix. To ensure a more accurate evaluation, we manually verified and corrected incorrect answers for the 1000 question-answer pairs, providing alternative correct answer candidates. The evaluation results on OCRBench are presented in Tab. 3, where Gemini achieves the highest score, followed by GPT4V in the second position. It is important to note that due to rigorous safety reviews by OPEN AI, GPT4V refused to provide results for 84 images in OCRBench. Monkey exhibited OCR capabilities that trailed only behind GPT4V and Gemini. From Tab. 3, we can observe that even state-of-the-art models like GPT4V and Gemini still struggle with the HMER task. Moreover, they also face challenges in processing unclear images, handwritten text, Non-Semantic text, and adhering to task instructions. As shown in Fig. 2 (g), even when explicitly requested to answer using the text found in the image, Gemini consistently interprets “02/02/2018” as “2 February 2018”.
@@
Why LMMs work for OCR? While we offer some analysis based on the results, the question as to why these multimodal models can deliver acceptable performance on OCR tasks remains challenging to conclusively explain. One plausible explanation for the success of multimodal models in Optical Character Recognition (OCR) tasks lies in the training data of multimodal models (similar to CLIP), which we believe includes some OCR data. However, unsupervised text-image pairs in the training data cannot compete with fully supervised data.
@@
From the perspective of architecture, the pre-trained visual encoder and Large Language Model (LLM) already demonstrate a solid understanding of their respective domain data, each working within their designated feature spaces. These LLMs connect visual and language data through elements like a linear projection layer, which acts as a visual tokenization step. By aligning visual tokens within the word embedding space of the pre-trained language model, visual embeddings closely resemble their corresponding text embeddings. This alignment facilitates text recognition, allowing the LLM to then present this OCR data to users in a generative manner. Future research could benefit from conducting an ablation analysis to better understand how the volume of multimodal training data impacts OCR performance.
Model Affiliation OCRBench Model Affiliation OCRBench
MiniCPM-V2.6 [62] OpenBMB 852 Cambrian-8B [63] NYU 614
InternVL2-Llama3-76B [64] Shanghai AI Lab 842 PaliGemma-3B-mix-448 [65] Google 614
CongRong [66] CloudWalk 827 Cambrian-13B [63] NYU 610
GLM-4v [67] Zhipu AI 814 MiniCPM-V-2 [68] OpenBMB 605
MiniMonkey-2B  [69] HUST 802 Cambrian-34B [63] NYU 591
InternVL2-8B [64] Shanghai AI Lab 794 CogVLM-17B-Chat [67] Zhipu AI 590
Claude3.5-Sonnet [70] Anthropic 788 LLaVA-Next-Yi-34B [71] UW–Madison 574
GPT-4o-mini-20240718 [72] OpenAI 785 TextMonkey [30] HUST 561
InternVL2-4B [64] Shanghai AI Lab 784 Monkey-Chat [29] HUST 534
InternVL2-2B [64] Shanghai AI Lab 781 InternLM-XComposer2  [73] Shanghai AI Lab 532
GLM-4v-9B [67] Zhipu AI 776 LLaVA-Next-Vicuna-7B [71] UW–Madison 532
CogVLM2-19B-Chat [67] Zhipu AI 757 LLaVA-Next-Mistral-7B [71] UW–Madison 531
InternVL2-1B [64] Shanghai AI Lab 755 RekaEdge [74] Reka AI 506
Gemini-1.5-Pro [75] Google 754 XVERSE-V-13B [76] XVERSE 489
Ovis1.5-Llama3-8B [77] Alibaba 744 Qwen-VL-Chat [78] Alibaba 488
Qwen-VL-Plus [78] Alibaba 726 InternLM-XComposer2-1.8B [73] Shanghai AI Lab 447
MiniCPM-Llama3-V2.5 [79] OpenBMB 725 Emu2_chat [80] BAAI 436
InternVL-Chat-V1.5 [64] Shanghai AI Lab 720 DeepSeek-VL-7B [81] DeepSeek 435
Claude3-Opus [82] Anthropic 694 OmniLMM-12B [83] OpenBMB 420
RekaFlash [74] Reka AI 692 TransCore-M [84] PCI Research 405
InternLM-XComposer2.5 [85] Shanghai AI Lab 686 LLaVA-InternLM2-7B [86] Shanghai AI Lab 402
Qwen-VL-Max [78] Alibaba 684 ShareGPT4V-13B [87] Shanghai AI Lab 398
Gemini-1.0-Pro [75] Google 680 360VL-70B [88] QiHoo360 397
InternLM-XComposer2-4KHD [89] Shanghai AI Lab 675 MiniCPM-V  [90] OpenBMB 366
Claude3-Haiku [82] Anthropic 658 Yi-VL-34B [91] 01-AI 290
Mini-InternVL-Chat-2B-V1.5 [64] Shanghai AI Lab 652 IDEFICS-80B-Instruct [92] Hugging Face 283
Claude3-Sonnet [82] Anthropic 646 LLaVA-Next-Llama3 [71] ByteDance 252
Mini-InternVL-Chat-4B-V1.5 [64] Shanghai AI Lab 639 MMAlaya [93] DataCanvas 223
Phi-3-Vision [94] Microsoft 637 VisualGLM [95] Tsinghua 170
WeMM [96] WechatCV 628 OpenFlamingo v2 [97] UW 149
IDEFICS2-8B [98] Hugging Face 626 PandaGPT-13B [99] Tencent AI Lab 36
Step-1V [100] StepFun 625 Chameleon-30B [101] Meta 27
Table 4: The results of OCRBench across various models are sourced from  [31] and their papers. Closed-source models are indicated in blue.
Future Directions: Although current LMMs have shown promising results, they still face challenges when dealing with complex tasks and are unable to match domain-specific methods in traditional text-related tasks. To enhance their performance, these models should concentrate on improving their ability to detect intricate features in images, enhancing their recognition of individual character shapes, and developing multilingual datasets for training.
@@
4 Future work
With the advancement of large multimodal models, numerous research institutions have made significant strides in optical character recognition (OCR) capabilities. For instance, MiniCPMV-2.6 achieved a score of 852 on OCRBench, while the Mini-Monkey scored 802 using just 2B parameters. These results underscore the immense potential of large multimodal models in the OCR domain, as highlighted in Table 4. OCRBench has been instrumental in advancing this field. However, it is important to note that the data utilized by most models is not entirely open-source, and many models remain proprietary. This demonstrates an existing opportunity for the open-source community to further explore and enhance the OCR capabilities of large multimodal models.
Furthermore, OCRBench currently lacks a comprehensive coverage of image data types and tasks. For instance, it falls short in encompassing more challenging data types like multilingual documents and texts captured in diverse scenarios, as well as tasks related to text detection. To address these gaps, our future work will focus on expanding OCRBench to include these elements, thereby fostering the ongoing advancement and utilization of large multimodal models in the realm of OCR.